OpenStudy (anonymous):

how do you solve 4th root of 16x^12 ?

OpenStudy (anonymous):

note that 16 = 2^4. and 12 = 4*3 so your expression becomes (2x^3)^4 the fourth root of this is?

jhonyy9 (jhonyy9):

\[2x^{3}\]